OpenStudy (anonymous):

How many customers does a portrait studio serve in a week if the studio averages 5 customers each weekday and averages twice as many customers on each weekend day?

OpenStudy (anonymous):

5*5 + 5*2*2 = 25 * 20 = 45 customers a week.

OpenStudy (anonymous):

5*5+10*2=45